# The Geometric View of Theories

### v1

@@ 2.3 Formulation of the models and dualities
By a ‘model’, I again mean a triple, $M=\langle{\cal S},{\cal Q},{\cal D}\rangle$, comprising a set of states $\cal S$}, i.e. a state-space, an algebra of quantities $\cal Q$} and a dynamics $\cal D$}, together with a rule for assigning values to quantities on states. (Additionally, the set of states and quantities usually comes equipped with symmetries, as automorphisms of these sets.) Because of this smooth assignment of values, the algebra of quantities and the state space are each other’s duals, in the mathematical, not the physical, sense. Thus, given a state $s\in{\cal S}$, a quantity $Q\in{\cal Q}$ assigns to it a (real or complex) value, i.e. $Q(s)=\langle Q,s\rangle$, where $\langle Q,\bullet\rangle$ is the rule for assigning values to states that is part of the definition of the model. Thus we have ${\cal Q}={\cal S}^{*}$, as well as ${\cal S}={\cal Q}^{*}$. For a duality, we take the state-space $\cal S$} to be the set of dynamically possible states, which are a subset of the kinematically possible states. [^25] Most classical dualities are valid at the level of the “off-shell” action, i.e. without using the equation of motion. Hence these dualities can be extended to all kinematical states. However, the element that is common to all dualities is that the dynamical states of duals are isomorphic, hence our restriction to those states. Thus, in effect, we can think of a model as a pair of a state-space and its dual, together with an assignment of a subspace of the state-space.
@@
Models in physics depend on pararameters such as coupling constants and expectation values of some select subset of quantities. This also follows from the ideas (1) and (2) in Section 1. When I wish to emphasise this, I will denote the model as: $M=M(t)$, where $t$ collectively denotes these parameters (see Section 3). As we discussed in Section 2.2, these parameters form a moduli space. [^27] Read (2016: pp. 228) discusses the idea of duals defined over a moduli space, and how they can be embedded in the parameter space of a successor theory. See also Rickles (2011: p. 65) and De Haro and Butterfield (2025: pp. 484-499).
@@ 3 A specific proposal: models in the fibres
This Section makes a structural proposal that (a) brings the remarks about quasi-dualities, moduli spaces, and fibre bundles from previous Sections, into a single structure; and (b) generalizes the examples of the geometric view that are discussed in two Appendices, and which aim to illustrate the main aspects of the view. (At this point, the interested reader is encouraged to read these Appendices.) I will propose that a physical theory is a fibre bundle constructed in a specific way, with states and quantities in the fibres, and a space of parameters or moduli in the base. I will call this a model bundle.
@@ 3.1.1 Definition of the model bundle and quasi-dualities
To define the model bundle, we first gather the threads from Section 2.3. Recall that a model is a triple $M=\langle{\cal S},{\cal Q},{\cal D}\rangle$ of structured set of states, structured set of quantities (usually, an algebra) and a dynamics. I will think of a model as a Cartesian product, $M={\cal Q}\times{\cal S}$. Thus my notation will suppress the dynamics, because the dynamics can be stated as a choice of a distinguished quantity in $\cal Q$} that is the Hamiltonian (or, alternatively, the Lagrangian). A duality then maps the Hamiltonian of one model to the Hamiltonian of the dual model. I will assume that this condition is satisfied, and will not need to indicate it in my notation.
@@
By analogy with our earlier examples, the obvious suggestion is to then put the models in the fibres. Thus we define a bundle $F\rightarrow E\overset{\pi}{\rightarrow}{\cal M}$, with total space $E$, fibres $F$, and projection $\pi:E\rightarrow{\cal M}$, i.e. ${\cal Q}(t)\times{\cal S}(t)\mapsto t\in{\cal M}$, that is a differentiable surjection, so that “nothing in the base manifold is left out”. If there are quasi-dualities, as I will in general assume, we put models that are related by a quasi-duality on the same fibre. As I will now explain, these quasi-dual models will be different elements of the fibre over a given point that are related by a diffeomorphism of the fibre. This is natural, because the quasi-dualities form a group that we can identify with the bundle’s structure group $G$, namely the group of fibre automorphisms, i.e. the diffeomorphisms of the fibre that preserve the structure of the fibre. This group acts vertically on the fibres and maps the models in the fibres to each other. More precisely, there are bundle charts, $\phi_{U}:E_{U}\rightarrow U\times F$, where $E_{U}\subset E$ and $U\subset{\cal M}$, that satisfy projection and consistency conditions. The consistency conditions, also called ‘cocycle conditions’, defined on overlaps of up to three regions, secure that the local pieces of the bundle can be glued consistently. Since the bundle charts allow us to write the bundle locally as a Cartesian product of the base and the fibre, they locally trivialize the bundle. By combining a chart and another chart’s inverse map, we can define transition functions $\phi_{j}\circ\phi_{i}^{-1}|_{(U_{i}\cap U_{j})\times F}:(U_{i}\cap U_{j})\times F\rightarrow(U_{i}\cap U_{j})\times F$. These transition functions are local diffeomorphisms of the bundle induced by the local mapping of the bundle into the Cartesian product. Note that, by the projection condition on the bundle charts, these transition functions act only vertically on the fibre, i.e. they leave the point $t\in U_{i}\cap U_{j}$ in the base fixed. This implies that we can use them to define a diffeomorphism that acts purely on the fibre above a point and leaves the base space fixed, i.e. its input and output are only the fibre. We do this by evaluating the transition function on a given point $t\in U_{i}\cap U_{j}$, and using a subscript in our notation: $\phi_{it}^{-1}:F\rightarrow E_{U_{i}}$, defined as $\phi^{-1}_{it}(f):=\phi_{i}^{-1}(t,f)$, where $(t,f)\in U_{i}\times F$. In this way we get, above every point $t\in U_{i}\cap U_{j}$ of the base, a diffeomorphism of the fibre: $\phi_{jt}\circ\phi_{it}^{-1}:F\rightarrow F$. These diffeomorphisms stay on a single fibre: they are elements of the structure group $G$ that acts vertically, which are the quasi-dualities. They contain the global information about the fibre: namely, they specify how the models in the fibre are glued. It is important to note that they are only defined above points of the base that lie in the intersection (overlap) of two regions of the kind $U$ that parametrize the set of charts, i.e. $t\in U_{i}\cap U_{j}\subset{\cal M}$. [^28] The fact that these diffeomorphisms are defined only locally, i.e. on overlapping regions, can also be emphasised as follows. The diffeomorphisms allow us to define maps from the basis into the structure group $G$ of diffeomorphisms of the fibre, $\phi_{ji}:U_{i}\cap U_{j}\rightarrow G$, given by $t\mapsto\phi_{jt}\circ\phi_{it}^{-1}$, where $t\in U_{i}\cap U_{j}$. These maps are also called transition functions (of the fibre). For an illustration of a model bundle, see Figure 1.
@@
The bundle charts defined above give us an interesting way to think about the elements of the fibre, i.e. the quasi-dual models $M_{i}(t)$, as different coordinatizations of the fibre $F_{t}$ over a point $t\in{\cal M}$. Each coordinatization is given by a transition function evaluated over the point $t$, i.e. $\phi_{it}$, which takes values in $F$. More precisely, the model $M_{i}(t)\subseteq F$ is the image of $\phi_{it}$, with as input all those points of the bundle, $e\in E_{U_{i}}$, whose projection is $t$, i.e. $\pi(e)=t\in U_{i}$. Thinking of models as coordinatizations of the fibre, which is itself coordinate-independent (as is the whole bundle), agrees with the contrast in De Haro and Butterfield (2025: p. 7) between a common core theory and its dual models. In that contrast, the models are representations (in the mathematical sense) of the common core theory, each model adding specific mathematical structure. Thus here, the fibre is the local common core (i.e. for a fixed value of the moduli) and its elements are models, which are coordinatizations of this local common core theory. This also justifies a way of speaking in physics, where it is often said that dualities and quasi-dualities map models that express the local common core theory in ‘different coordinates’. This is literally the case for a model bundle, because the local trivializations of the bundle give us choices of coordinates where the quasi-dualities between models are changes of coordinates. [^31] As I emphasised in Section 2.2-(2), this way of speaking uses the additional differential structure of the fibre bundle, as against a bundle or a sheaf. In other examples, algebraic rather than differential structure may be required: see footnote 34.
@@ 3.1.2 Recovering dualities
This can also be understood in the opposite direction: a map between models can fall short of being a duality, and thus be a quasi-duality, in two ways (see the beginning of Section 2.2): through its failing to be a bijection, or through its failing to preseve a common core that is a theory. We will briefly discuss non-invertible quasi-dualities, i.e. quasi-dualities that fail to be bijections, in the next Section. The failure of the quasi-dualities to extend to the whole moduli space, that we discussed above, is indeed a case of the quasi-duality’s not preserving a common core theory for all values of $t$: in other words, the quasi-duality is defined only for fixed or special values of $t$, but not throughout the whole moduli space. (Recall, from above, that the fact that some quasi-dualities relate models at different values of the coupling is consistent with the fact that both dualities and quasi-dualities act locally, on the fibre above a point. For the illustration of this fact in the Seiberg-Witten theory, see footnote 30.)
@@ 3.1.3 Generalizations of the construction
The model bundle’s structure can be weakened in two obvious directions: ‘vertically’ and ‘horizontally’. (i) In the vertical direction, we can weaken the action of the structure group $G$ by taking it not to be a diffeomorphism of the fibre. This means that the quasi-dualities are not local isomorphisms, i.e. isomorphisms on the local fibres (recall that, in a model bundle, a quasi-duality is taken to be a local diffeomorphism). As usual, this can happen in either of two ways: (1) Through the quasi-duality’s not being structure-preserving: this is the case in the Seiberg-Witten theory, where we simplified the theory by not taking into account its spin structure. Taking into account this spin structure would give a bijection that is not locally structure-preserving. (2) Through the quasi-duality’s not being a local bijection. For example, the quasi-duality might be non-invertible, and then one would get a semi-group, rather than a group, $G$. Non-invertible quasi-dualities appear in recent discussions of the quantum Ising model. [^33] A prominent example is non-invertible Kramers-Wannier quantum quasi-duality: see Seiberg and Shao (2023) and Shao (2023). Combining non-invertible quasi-dualities with a bundle structure results in a semi-group model bundle. I am not aware of major uses of these structures in physics. (ii) In the horizontal direction, we can weaken the assumption that there is a generic fibre $F$ to which the fibres over each point $t$ are isomorphic. Thus one might consider a bundle that is not a fibre bundle (see e.g. Isham, 1999: Section 5.1) or, more generally a sheaf. [^34] One of the physical motivations for considering sheaves is the need to define vector bundles over submanifolds of a space (Sharpe 2003: Section 2.3). At its simplest, a sheaf associates a mathematical object, such as a set, a group, a module or a category, to every open set on a topological space. Informally, the definition of a sheaf usually begins by associating sections, i.e. the mathematical objects under consideration, to the open sets of the topological space, and then considering restriction maps on these sections under nesting of the open sets. To obtain a sheaf, additional conditions are required; see Eisenbund and Harris (2000: pp. 11-12). In quantum field theory and string theory, certain D-brane states on Calabi-Yau manifolds are related to sheaves over a moduli space, which can be used to understand mirror symmetry, as an instance of the (geometric) Langlands correspondence. More precisely, the geometric Langlands correspondence is realized as an equivalence between: (i) certain electric branes, defined on the moduli space ${\cal M}(G,C)$ of flat holomorphic $G$-bundles on a compact Riemann surface $C$, which are solutions of the Hitchin equations with gauge group $G$ on $C$; and (ii) magnetic branes, which are described by a sheaf of differential operators on the moduli space with Langlands-dual group $^{L}G$ (for a simple definition of the Langlands-dual group, see De Haro and Butterfield, 2025, p. 243; Gukov and Witten 2006, pp. 168–175). The moduli space ${\cal M}(G,C)$ has singularities whose physical interpretation is similar to that in the Seiberg-Witten theory. In the neighbourhood of these singularities, the low-energy sigma models that are used to realize the Langlands correspondence are not valid, and one has to go “up” to four dimensions, to a version of ${\cal N}=4$ supersymmetric Yang-Mills theory. The Langlands correspondence originates in the $S$-duality of this ${\cal N}=4$ supersymmetric Yang-Mills theory. For discussions, see Kapustin and Witten (2007) and Kapustin (2008), who focus on the geometric Langlands programme; and Gaiotto and Witten (2022), for its analytic version.

### v2

@@ 2.3 Formulation of the models and dualities
By a ‘model’, I again mean a triple, $M=\langle{\cal S},{\cal Q},{\cal D}\rangle$, comprising a set of states $\cal S$}, i.e. a state-space, an algebra of quantities $\cal Q$} and a dynamics $\cal D$}, together with a rule for assigning values to quantities on states. (Additionally, the set of states and quantities usually comes equipped with symmetries, as automorphisms of these sets.) Because of this smooth assignment of values, the algebra of quantities and the state space are each other’s duals, in the mathematical, not the physical, sense. Thus, given a state $s\in{\cal S}$, a quantity $Q\in{\cal Q}$ assigns to it a (real or complex) value, i.e. $Q(s)=\langle Q,s\rangle$, where $\langle Q,\bullet\rangle$ is the rule for assigning values to states that is part of the definition of the model. Thus we have ${\cal Q}={\cal S}^{*}$, as well as ${\cal S}={\cal Q}^{*}$. For a duality, we take the state-space $\cal S$} to be the set of dynamically possible states, which are a subset of the kinematically possible states. [^29] Most classical dualities are valid at the level of the “off-shell” action, i.e. without using the equations of motion. Hence these dualities can be extended to all kinematical states. However, the element that is common to all dualities is that the dynamical states of duals are isomorphic, hence our restriction to those states. Thus, in effect, we can think of a model as a pair of a state-space and its dual, together with an assignment of a subspace of the state-space.
@@
Models in physics depend on pararameters such as coupling constants and expectation values of some select subset of quantities. This also follows from the ideas (1) and (2) in Section 1. When I wish to emphasise this, I will denote the model as: $M=M(t)$, where $t$ collectively denotes these parameters (see Section 3). As we discussed in Section 2.2, these parameters form a moduli space. [^31] Read (2016: pp. 228) discusses the idea of duals defined over a moduli space, and how they can be embedded in the parameter space of a successor theory. See also Rickles (2011: p. 65) and De Haro and Butterfield (2025: pp. 484–499).
@@ 3 A specific proposal: models in the fibres
This Section makes a structural proposal that (a) brings the remarks about quasi-dualities, moduli spaces, and fibre bundles from previous Sections, into a single structure; and (b) generalizes the examples of the geometric view that are discussed in two Appendices, and which aim to illustrate the main aspects of the view. (At this point, the interested reader is encouraged to read these Appendices.) I will propose that, at its simplest, a physical theory is a fibre bundle constructed in a specific way, with states and quantities in the fibres, and a space of parameters or moduli in the base. I will call this a model bundle.
@@ 3.1.1 Definition of the model bundle and quasi-dualities
To define the model bundle, we first gather the threads from Section 2.3. Recall that a model is a triple $M=\langle{\cal S},{\cal Q},{\cal D}\rangle$ of structured set of states, structured set of quantities (usually, an algebra) and a dynamics. I will think of a model as a Cartesian product, $M={\cal Q}\times{\cal S}$. Thus my notation will suppress the dynamics, because in most cases of interest the dynamics can be stated as a choice of a distinguished quantity in $\cal Q$} that is the Hamiltonian (or, alternatively, the Lagrangian). A duality then maps the Hamiltonian of one model to the Hamiltonian of the dual model. I will assume that this condition is satisfied, and will not need to indicate it in my notation. [^32] De Haro and Butterfield (2025: p. 72) require that the duality map is equivariant for the two triples’ dynamics (in either Schrödinger or Heisenberg pictures). For systems described by a Hamiltonian, this condition translates into the condition that the Hamiltonians match under the duality map. For example, in a quantum theory in the Schrödinger picture, the time evolution is implemented by the action of a unitary operator $U$ on states. The equivariance condition then comes down to the requirement that the time evolution operators of the two duals are related by the duality map $d_{\cal S}$ on states, as follows: $U_{2}=d_{\cal S}\,U_{1}\,d_{\cal S}^{-1}$.
@@
By analogy with our earlier examples, the obvious suggestion is to then put the models in the fibres. [^33] In the context of inter-theoretic reduction, Landsman (2017), Feintzeig (2020, 2022), and Steeger and Feintzeig (2021) define a continuous bundle of C*-algebras over a locally compact Hausdorff space, where the fibres are a family of C*-algebras, the base is a parameter space, and the continuous sections assign to each point in the base an element of the corresponding C*-algebra. Here, the parameter space is typically the set of values of Planck’s constant $\hbar$ (in the case of quantum mechanics) and, in cases involving a system with $N$ degrees of freedom (with $N$ large), $\{1/N|N\in\mathbb{N}\}\cup\{0\}$. This illustrates how the idea of a model bundle can also be used to discuss inter-theoretic relations like reduction: in these works, the bundle construction is used to investigate the quantum-to-classical transition (i.e. sending $\hbar\rightarrow 0$) or a transition between a system with a finite and with an infinite number of degrees of freedom (e.g. a quantum spin system, sending $N\rightarrow\infty$). Note that a continuous bundle of C*-algebras is not a fibre bundle, because no local trivialization exists at $\hbar=0$. See Landsman (2017: pp. 294, 297) and Feintzeig (2020: p. 621).
@@
Thus we define a bundle $F\rightarrow E\overset{\pi}{\rightarrow}{\cal M}$, with total space $E$, fibres $F$, and projection $\pi:E\rightarrow{\cal M}$, i.e. ${\cal Q}(t)\times{\cal S}(t)\mapsto t\in{\cal M}$, that is a differentiable surjection, so that “nothing in the base manifold is left out”. If there are quasi-dualities, as I will in general assume, we put models that are related by a quasi-duality on the same fibre. As I will now explain, these quasi-dual models will be different elements of the fibre over a given point that are related by a diffeomorphism of the fibre. This is natural, because the quasi-dualities form a group that we can identify with the bundle’s structure group $G$, namely the group of fibre automorphisms, i.e. the diffeomorphisms of the fibre that preserve the structure of the fibre. This group acts vertically on the fibres and maps the models in the fibres to each other. More precisely, there are bundle charts, $\phi_{U}:E_{U}\rightarrow U\times F$, where $E_{U}\subset E$ and $U\subset{\cal M}$, that satisfy projection and consistency conditions. The consistency conditions, also called ‘cocycle conditions’, defined on overlaps of up to three regions, secure that the local pieces of the bundle can be glued consistently. Since the bundle charts allow us to write the bundle locally as a Cartesian product of the base and the fibre, they locally trivialize the bundle. By combining a chart and another chart’s inverse map, we can define transition functions $\phi_{j}\circ\phi_{i}^{-1}|_{(U_{i}\cap U_{j})\times F}:(U_{i}\cap U_{j})\times F\rightarrow(U_{i}\cap U_{j})\times F$. These transition functions are local diffeomorphisms of the bundle induced by the local mapping of the bundle into the Cartesian product. Note that, by the projection condition on the bundle charts, these transition functions act only vertically on the fibre, i.e. they leave the point $t\in U_{i}\cap U_{j}$ in the base fixed. This implies that we can use them to define a diffeomorphism that acts purely on the fibre above a point and leaves the base space fixed, i.e. its input and output are only the fibre. We do this by evaluating the transition function on a given point $t\in U_{i}\cap U_{j}$, and using a subscript in our notation: $\phi_{it}^{-1}:F\rightarrow E_{U_{i}}$, defined as $\phi^{-1}_{it}(f):=\phi_{i}^{-1}(t,f)$, where $(t,f)\in U_{i}\times F$. In this way we get, above every point $t\in U_{i}\cap U_{j}$ of the base, a diffeomorphism of the fibre: $\phi_{jt}\circ\phi_{it}^{-1}:F\rightarrow F$. These diffeomorphisms stay on a single fibre: they are elements of the structure group $G$ that acts vertically, which are the quasi-dualities. They contain the global information about the fibre: namely, they specify how the models in the fibre are glued. It is important to note that they are only defined above points of the base that lie in the intersection (overlap) of two regions of the kind $U$ that parametrize the set of charts, i.e. $t\in U_{i}\cap U_{j}\subset{\cal M}$. [^34] The fact that these diffeomorphisms are defined only locally, i.e. on overlapping regions, can also be emphasised as follows. The diffeomorphisms allow us to define maps from the basis into the structure group $G$ of diffeomorphisms of the fibre, $\phi_{ji}:U_{i}\cap U_{j}\rightarrow G$, given by $t\mapsto\phi_{jt}\circ\phi_{it}^{-1}$, where $t\in U_{i}\cap U_{j}$. These maps are also called transition functions (of the fibre). For an illustration of a model bundle, see Figure 1.
@@
The bundle charts defined above give us an interesting way to think about the elements of the fibre, i.e. the quasi-dual models $M_{i}(t)$, as different coordinatizations of the fibre $F_{t}$ over a point $t\in{\cal M}$. Each coordinatization is given by a transition function evaluated over the point $t$, i.e. $\phi_{it}$, which takes values in $F$. More precisely, the model $M_{i}(t)\subseteq F$ is the image of $\phi_{it}$, with as input all those points of the bundle, $e\in E_{U_{i}}$, whose projection is $t$, i.e. $\pi(e)=t\in U_{i}$. Thinking of models as coordinatizations of the fibre, which is itself coordinate-independent (as is the whole bundle), agrees with the contrast in De Haro and Butterfield (2025: p. 7) between a common core theory and its dual models. In that contrast, the models are representations (in the mathematical sense) of the common core theory, each model adding specific mathematical structure. Thus here, the fibre is the local common core (i.e. for a fixed value of the moduli) and its elements are models, which are coordinatizations of this local common core theory. This also justifies a way of speaking in physics, where it is often said that dualities and quasi-dualities map models that express the local common core theory in ‘different coordinates’. This is literally the case for a model bundle, because the local trivializations of the bundle give us choices of coordinates where the quasi-dualities between models are changes of coordinates. [^37] As I emphasised in Section 2.2-(2), this way of speaking uses the additional differential structure of the fibre bundle, as against a bundle or a sheaf. In other examples, algebraic rather than differential structure may be required: see footnote 40.
@@ 3.1.2 Recovering dualities
This can also be understood in the opposite direction: a map between models can fall short of being a duality, and thus be a quasi-duality, in two ways (see the beginning of Section 2.2): through its failing to be a bijection, or through its failing to preseve a common core that is a theory. We will briefly discuss non-invertible quasi-dualities, i.e. quasi-dualities that fail to be bijections, in the next Section. The failure of the quasi-dualities to extend to the whole moduli space, that we discussed above, is indeed a case of the quasi-duality’s not preserving a common core theory for all values of $t$: in other words, the quasi-duality is defined only for fixed or special values of $t$, but not throughout the whole moduli space. (Recall, from above, that the fact that some quasi-dualities relate models at different values of the coupling is consistent with the fact that both dualities and quasi-dualities act locally, on the fibre above a point. For the illustration of this fact in the Seiberg-Witten theory, see footnote 36.)
@@ 3.1.3 Generalizations of the construction
The model bundle’s structure can be weakened in two obvious directions: ‘vertically’ and ‘horizontally’. (i) In the vertical direction, we can weaken the action of the structure group $G$ by taking it not to be a diffeomorphism of the fibre. This means that the quasi-dualities are not local isomorphisms, i.e. isomorphisms on the local fibres (recall that, in a model bundle, a quasi-duality is taken to be a local diffeomorphism). As usual, this can happen in either of two ways: (1) Through the quasi-duality’s not being structure-preserving: this is the case in the Seiberg-Witten theory, where we simplified the theory by not taking into account its spin structure. Taking into account this spin structure would give a bijection that is not locally structure-preserving. (2) Through the quasi-duality’s not being a local bijection. For example, the quasi-duality might be non-invertible, and then one would get a semi-group, rather than a group, $G$. Non-invertible quasi-dualities appear in recent discussions of the quantum Ising model. [^39] A prominent example is non-invertible Kramers-Wannier quantum quasi-duality: see Seiberg and Shao (2023) and Shao (2023). Combining non-invertible quasi-dualities with a bundle structure results in a semi-group model bundle. I am not aware of major uses of these structures in physics. (ii) In the horizontal direction, we can weaken the assumption that there is a generic fibre $F$ to which the fibres over each point $t$ are isomorphic. Thus one might consider a bundle that is not a fibre bundle (see e.g. Isham, 1999: Section 5.1) or, more generally a sheaf. [^40] One of the physical motivations for considering sheaves is the need to define vector bundles over submanifolds of a space (Sharpe 2003: Section 2.3). At its simplest, a sheaf associates a mathematical object, such as a set, a group, a module or a category, to every open set on a topological space. Informally, the definition of a sheaf usually begins by associating sections, i.e. the mathematical objects under consideration, to the open sets of the topological space, and then considering restriction maps on these sections under nesting of the open sets. To obtain a sheaf, additional conditions are required; see Eisenbund and Harris (2000: pp. 11–12). In quantum field theory and string theory, certain D-brane states on Calabi-Yau manifolds are related to sheaves over a moduli space, which can be used to understand mirror symmetry, as an instance of the (geometric) Langlands correspondence. More precisely, the geometric Langlands correspondence is realized as an equivalence between: (i) certain electric branes, defined on the moduli space ${\cal M}(G,C)$ of flat holomorphic $G$-bundles on a compact Riemann surface $C$, which are solutions of the Hitchin equations with gauge group $G$ on $C$; and (ii) magnetic branes, which are described by a sheaf of differential operators on the moduli space with Langlands-dual group $^{L}G$ (for a simple definition of the Langlands-dual group, see De Haro and Butterfield, 2025, p. 243; Gukov and Witten 2006, pp. 168–175). The moduli space ${\cal M}(G,C)$ has singularities whose physical interpretation is similar to that in the Seiberg-Witten theory. In the neighbourhood of these singularities, the low-energy sigma models that are used to realize the Langlands correspondence are not valid, and one has to go “up” to four dimensions, to a version of ${\cal N}=4$ supersymmetric Yang-Mills theory. The Langlands correspondence originates in the $S$-duality of this ${\cal N}=4$ supersymmetric Yang-Mills theory. For discussions, see Kapustin and Witten (2007) and Kapustin (2008), who focus on the geometric Langlands programme; and Gaiotto and Witten (2022), for its analytic version.